# Towards Multi-Modal Animal Pose Estimation: An In-Depth Analysis

### v1

@@ A. Extensions of LEAP method
A notable extension of the LEAP (Pereira et al., 2019) method is T-LEAP (Russello et al., 2022), which conceptualises temporal information as an auxiliary dimension within the CNN framework, thereby employing 3D convolutional layers to capture both spatial and temporal features concurrently to extract animal poses from video data. To leverage the temporal information embedded in videos, T-LEAP processes a sequence of frames of length $T$, and outputs the confidence maps linked to the last frame $t=T$. As for architectural modifications compared to a deeper version of LEAP, all convolution, pooling and transposed-convolution operations transition from 2D to 3D. Notably, these 3D convolutional operations are extended to not only the spatial dimension of the input but also to the temporal dimension, thus facilitating the extraction of spatiotemporal signals from the video data, and making T-LEAP robust for occlusions. For example, in a scenario where a cow’s leg is discernible in one frame but obscured in the subsequent frame, T-LEAP can employ visible information to predict the probable position of the leg in the occluded frame. Another extension of the LEAP model is SLEAP (Pereira et al., 2022), which also incorporates temporal information, but unlike LEAP, this extends the model to multi-animal pose estimation.
B. Optical Flow-based Methods
@@
OptiFlex (Liu et al., 2021) is an optical flow-based methods that allow the estimation of motion vectors between consecutive frames to capture the temporal dynamics of animal poses. In particular, OptiFlex (Liu et al., 2021) combines a deep learning-based pose estimation module called FlexibleBaseline with an optical flow module named OpticalFlow to enhance its APE performance across multiple frames. The FlexibleBaseline module predicts the initial pose for each frame, and the OpticalFlow module converges initial predictions on a target frame and its adjacent frames into a final prediction for the target frame. By tracking the pixel-level displacements over time, optical flow estimation techniques, such as the Lucas-Kanade algorithms (Baker and Matthews, 2004), can compute dense motion fields between frames, thereby inferring the movement of body parts and joints. Suppose a certain body part is occluded or unclear in a frame. In that case, the model can infer its likely position based on the detected motion and poses in adjacent frames, showing its robustness against temporary obstruction of keypoint poses.
C. Future Trend using Semi-supervised and Unsupervised Video Adaptation Learning
@@
Recently, there has been a growing shift towards using semi-supervised learning (e.g. SemiMultiPose (Blau et al., 2022), Lightning Pose (Biderman et al., 2023)) and unsupervised video adaptation learning (e.g. SuperAnimal (Shaokai Ye and Mathis, 2024)) to extract poses from videos. In 2D single-frame APE methods, supervised learning dominates, relying on many labelled frames for training. However, manually adding annotations for videos is costly and labour-intensive, especially for multiple instances. In real-world scenarios like wildlife videography, there are exponentially more abundant unlabeled frames than labelled frames. Given these deficiencies, SemiMultiPose (Blau et al., 2022) proposes a semi-supervised architecture to exploit the abundant structures pervasive in these unlabeled video frames to enhance the training process, which is particularly useful in the sparse labelling video-based scenarios. Similarly, Lightning Pose (2023) (Biderman et al., 2023) incorporates a semi-supervised approach to exploit the spatiotemporal statistics present in unlabeled videos. This methodology introduces unsupervised training objectives that penalise the network when its predictions deviate from the expected smoothness of physical motion, multiple-camera-views geometry, or stray from a low-dimensional subspace of plausible body configurations. Additionally, Lightning Pose predicts the pose for a specific frame by considering the temporal context from surrounding unlabeled frames. This incorporation of context is instrumental in resolving transient occlusions and ambiguities between proximal and morphologically similar body parts. Last, the SuperAnimal (2024) (Shaokai Ye and Mathis, 2024), an unsupervised video adaptation APE method, can be applied to images and videos. However, its approach differs depending on the data type. For images, SuperAnimal operates as a supervised method, relying on labelled data for training. In contrast, for videos, it functions as an unsupervised method by incorporating unsupervised techniques, such as video adaptation and pseudo-labelling, to improve performance on unseen videos. This method helps reduce temporal jitter and refine model outputs without further manual annotation. However, the unsupervised techniques are applied after the model has been initially trained in a supervised manner. These works demonstrate better APE performance with fewer labels, generalise better to unseen videos, and provide smoother and more reliable pose trajectories for downstream analysis, which proves the shift towards semi-supervised and unsupervised learning in APE.
4.3. 3D Animal Pose Estimation from Images and Videos
@@
3D APE refers to the task of determining the locations of animal body joints in 3D space. This challenge extends beyond the 2D APE by adding depth information to the analysis, which allows for a deeper understanding of animal behaviour, biomechanics, and motion. Some 3D APE methods extract poses from monocular imagery and video data streams. However, these singular perspectives introduce an ill-posed nature to the problem, as the reduction from 3D reality to 2D representation inherently results in the loss of spatial data. Conversely, the integration of multiple-camera-views imaging facilitates a well-conditioned approach to interpolate the lost dimension for 3D APE using geometric consistency. Based on the number of camera views used to extract the animal poses, we categorize the 3D APE into two sections: (1) 3D APE from single-camera-view images and videos (Salem et al., 2019; Gosztolai et al., 2021; Dai et al., 2023; Sosa and Hogg, 2023; Shooter et al., 2024; Hu et al., 2023; Zuffi et al., 2017, 2019a; Biggs et al., 2020; Li and Lee, 2021a; Wang et al., 2021a; Rüegg et al., 2022, 2023; Wu et al., 2023a, b; Yang et al., 2022a; Li et al., 2024b), (2) 3D APE from multiple-camera-views images and videos (Yao et al., 2019; Bala et al., 2020; Zhang et al., 2021; Karashchuk et al., 2021; Dunn et al., 2021; Muramatsu et al., 2022; Ebrahimi et al., 2023; An et al., 2023; Han et al., 2024).
@@
Table 7. 3D RGB-based APE methods. All the symbols and column meanings are defined in table 5 and table 6, and any additional symbols are defined here: [Uncaptioned image] is shape model (Zuffi et al., 2017), [Uncaptioned image] is Structured Forests architecture (Dollár and Zitnick, 2015), [Uncaptioned image] is Graph Convolutional Network (GCN), [Uncaptioned image] is Stacked Hourglass architecture based on CNN, [Uncaptioned image] is GAN, [Uncaptioned image] is the combination of Neural Radiance Fields (NeRF), Volumetric Signed Distance Functions (VolSDF) & Differentiable Physics Simulation, [Uncaptioned image] is a framework based on Epipolar Divergence, [Uncaptioned image] is Gaussian-based model, [Uncaptioned image] is Statistical Shape Model (SSM), [Uncaptioned image] is NeRF, [Uncaptioned image] is Semantic Bank of Skinned Models (SBSM), [Uncaptioned image] is feed-forward neural network. [Uncaptioned image] is the Body Mesh Recovery method
@@
3D APE from Image & Video - Single-Camera-View Methods:
@@
1 2019 Salem et al. (Salem et al., 2019) I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mouse Pose Distance Measure (Perona, 2010),
@@
2 2021 LiftPose3D (Gosztolai et al., 2021) V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mice, Rats & Macaques MAE
3 2023 Hu et al. (Hu et al., 2023) V S {[Uncaptioned image], [Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mouse OKS for 2D Keypoints,
@@
4 2023 Dai et al. (Dai et al., 2023) I U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AcinoSet (Joska et al., 2021), N-MPJPE
SMAL (Zuffi et al., 2017)
5 2023 Sosa et al. (Sosa and Hogg, 2023) I SF [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Horse PCK@0.05 for 2D,
(Weizmann Data) (Borenstein et al., 2004) Qualitative for 3D
6 2024 DigiDogs/D-Pose (Shooter et al., 2024) I SF [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Dogs PCK for 2D&3D, MPJPE for 3D
AMR (Model-Based):
7 2017 SMAL (Zuffi et al., 2017) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} Quadruped Scan-to-Mesh Distance,
Keypoint Reprojection Err.,
Silhouette Reprojection Err.
8 2019 Three-D Safari (Zuffi et al., 2019a) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Zebra PCK, IoU, Geodesic Distance
9 2020 SMBLD (Biggs et al., 2020) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} StanfordExtra (Biggs et al., 2020) IoU, PCK
10 2021 Li et al. (Li and Lee, 2021a) I S [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]} [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} StanfordExtra (Biggs et al., 2020), PCK, IoU
Animal Pose (Cao et al., 2019),
BADJA (Biggs et al., 2018)
11 2021 Birds of a Feather (Wang et al., 2021a) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Birds, StanfordExtra (Biggs et al., 2020) PCK, IoU
12 2022 BARC (Rüegg et al., 2022) I S [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} StanfordExtra (Biggs et al., 2020) PCK & IoU for 2D,
@@
13 2023 BITE (Rüegg et al., 2023) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} StanfordExtra (Biggs et al., 2020) Scan-to-Mesh Distance,
Semi-Synthetic 3D (Rüegg et al., 2023) Mesh-to-Scan Distance,
@@
AMR (Model-Free):
14 2023 DOVE (Wu et al., 2023a) I,V U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Bird, Horse, mIoU,
@@
15 2023 MagicPony (Wu et al., 2023b) I U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Horse-10 (Mathis et al., 2021), PCK,
DOVE Data (Wu et al., 2023a), 3D Chamfer Distance (cm)
COCO (Lin et al., 2014a), & others
16 2022 BANMo (Yang et al., 2022a) I,V SF [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]} , [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} Animals & Humans 3D Chamfer Distance (cm),
F-score
17 2024 3D-Fauna (Li et al., 2024b) I U [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image] Quadruped Animals PCK@0.1, KT-PCK@0.1
3D APE from Image & Video - Multi-Camera-Views Methods:
18 2019 Monet (Yao et al., 2019) I SM [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Dog, Monkey PCK, Reprojection Err.
19 2020 OpenMonkeyStudio (Bala et al., 2020) I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Monkey Median & Mean Err. (cm),
Standard Deviation (cm)
20 2021 GIMBAL (Zhang et al., 2021) V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mouse MPE (mm), RPA-MPE (mm)
21 2021 DAN-NCE (Dunn et al., 2021) V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Rats, Marmosets Mice PCK, Reconstruction Err.
@@
22 2021 Anipose (Karashchuk et al., 2021) V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Fly, Mouse Joint Position Err. (mm)
23 2022 Muramatsu et al. (Muramatsu et al., 2022) V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AcinoSet (Joska et al., 2021) RMSE, PCK,
$l^{2}$ norm in pixels
24 2023 3D-UPPER (Ebrahimi et al., 2023) V U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mice RMSE, Scalar Product,
@@
25 2024 MAMMAL (An et al., 2023) V S {[Uncaptioned image], [Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Pigs, Mouse, Dogs MPJPE, IoU,
@@
26 2024 SBeA (Han et al., 2024) V SF [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image] Mice, Dogs, Birds RMSE (pixels), Reprojection Err.
4.3.1. Single Camera View

### v2

@@ 4.2.2 End-to-End Methods to Incorporate Temporal Information
Extensions of LEAP Method: A notable extension of the LEAP [111] method is T-LEAP [122], which conceptualises temporal information as an auxiliary dimension within the CNN framework, thereby employing 3D convolutional layers to capture both spatial and temporal features concurrently to extract animal poses from video data. To leverage the temporal information embedded in videos, T-LEAP processes a sequence of frames of length $T$, and outputs the confidence maps linked to the last frame $t=T$. As for architectural modifications compared to a deeper version of LEAP, all convolution, pooling and transposed-convolution operations transition from 2D to 3D. Notably, these 3D convolutional operations are extended to not only the spatial dimension of the input but also to the temporal dimension, thus facilitating the extraction of spatiotemporal signals from the video data, and making T-LEAP robust for occlusions. For example, in a scenario where a cow’s leg is discernible in one frame but obscured in the subsequent frame, T-LEAP can employ visible information to predict the probable position of the leg in the occluded frame. Another extension of the LEAP model is SLEAP [112], which also incorporates temporal information, but unlike LEAP, this extends the model to multi-animal pose estimation.
@@
Optical Flow-based Methods: OptiFlex [81] is an optical flow-based methods that allow the estimation of motion vectors between consecutive frames to capture the temporal dynamics of animal poses. In particular, OptiFlex [81] combines a deep learning-based pose estimation module called FlexibleBaseline with an optical flow module named OpticalFlow to enhance its APE performance across multiple frames. The FlexibleBaseline module predicts the initial pose for each frame, and the OpticalFlow module converges initial predictions on a target frame and its adjacent frames into a final prediction for the target frame. By tracking the pixel-level displacements over time, optical flow estimation techniques, such as the Lucas-Kanade algorithms [7], can compute dense motion fields between frames, thereby inferring the movement of body parts and joints. Suppose a certain body part is occluded or unclear in a frame. In that case, the model can infer its likely position based on the detected motion and poses in adjacent frames, showing its robustness against temporary obstruction of keypoint poses.
@@
Future Trend using Semi-supervised and Unsupervised Video Adaptation Learning: Recently, there has been a growing shift towards using semi-supervised learning (e.g. SemiMultiPose [12], Lightning Pose [9]) and unsupervised video adaptation learning (e.g. SuperAnimal [160]) to extract poses from videos. In 2D single-frame APE methods, supervised learning dominates, relying on many labelled frames for training. However, manually adding annotations for videos is costly and labour-intensive, especially for multiple instances. In real-world scenarios like wildlife videography, there are exponentially more abundant unlabeled frames than labelled frames. Given these deficiencies, SemiMultiPose [12] proposes a semi-supervised architecture to exploit the abundant structures pervasive in these unlabeled video frames to enhance the training process, which is particularly useful in the sparse labelling video-based scenarios. Similarly, Lightning Pose (2023) [9] incorporates a semi-supervised approach to exploit the spatiotemporal statistics present in unlabeled videos. This methodology introduces unsupervised training objectives that penalise the network when its predictions deviate from the expected smoothness of physical motion, multiple-camera-views geometry, or stray from a low-dimensional subspace of plausible body configurations. Additionally, Lightning Pose predicts the pose for a specific frame by considering the temporal context from surrounding unlabeled frames. This incorporation of context is instrumental in resolving transient occlusions and ambiguities between proximal and morphologically similar body parts. Last, the SuperAnimal (2024) [160], an unsupervised video adaptation APE method, can be applied to images and videos. However, its approach differs depending on the data type. For images, SuperAnimal operates as a supervised method, relying on labelled data for training. In contrast, for videos, it functions as an unsupervised method by incorporating unsupervised techniques, such as video adaptation and pseudo-labelling, to improve performance on unseen videos. This method helps reduce temporal jitter and refine model outputs without further manual annotation. However, the unsupervised techniques are applied after the model has been initially trained in a supervised manner. These works demonstrate better APE performance with fewer labels, generalise better to unseen videos, and provide smoother and more reliable pose trajectories for downstream analysis, which proves the shift towards semi-supervised and unsupervised learning in APE.
4.3 3D Animal Pose Estimation from Images and Videos
@@
3D APE refers to the task of determining the locations of animal body joints in 3D space. This challenge extends beyond the 2D APE by adding depth information to the analysis, which allows for a deeper understanding of animal behaviour, biomechanics, and motion. Some 3D APE methods extract poses from monocular imagery and video data streams. However, these singular perspectives introduce an ill-posed nature to the problem, as the reduction from 3D reality to 2D representation inherently results in the loss of spatial data. Conversely, the integration of multiple-camera-views imaging facilitates a well-conditioned approach to interpolate the lost dimension for 3D APE using geometric consistency. Based on the number of camera views used to extract the animal poses, we categorize the 3D APE into two sections: (1) 3D APE from single-camera-view images and videos [123, 39, 19, 132, 128, 47, 178, 176, 10, 68, 144, 121, 147, 120, 148, 153, 74], (2) 3D APE from multiple-camera-views images and videos [159, 8, 168, 55, 30, 96, 31, 2, 43].
@@
Table 7: 3D RGB-based APE methods. All the symbols and column meanings are defined in Table 5 and Table 6, and any additional symbols are defined here: [Uncaptioned image] is shape model [178], [Uncaptioned image] is Structured Forests architecture [27], [Uncaptioned image] is Graph Convolutional Network (GCN), [Uncaptioned image] is Stacked Hourglass architecture based on CNN, [Uncaptioned image] is GAN, [Uncaptioned image] is the combination of Neural Radiance Fields (NeRF), Volumetric Signed Distance Functions (VolSDF) & Differentiable Physics Simulation, [Uncaptioned image] is a framework based on Epipolar Divergence, [Uncaptioned image] is Gaussian-based model, [Uncaptioned image] is Statistical Shape Model (SSM), [Uncaptioned image] is NeRF, [Uncaptioned image] is Semantic Bank of Skinned Models (SBSM), [Uncaptioned image] is feed-forward neural network. [Uncaptioned image] is the Body Mesh Recovery method. The method citation is available in column Year.
@@
3D APE from Image & Video - Single-Camera-View Methods
@@
1 2019 Salem et al. I,V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mouse Pose Distance Measure (2010),
@@
2 2021 LiftPose3D V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mice, Rats, Macaques MAE
3 2023 Hu et al. V S {[Uncaptioned image], [Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mouse OKS for 2D Keypoints,
@@
4 2023 Dai et al. I U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] SMAL (2017), N-MPJPE
AcinoSet (2021)
5 2023 Sosa et al. I SF [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Horse - Weizmann (2004) PCK@0.05 for 2D,
Qualitative for 3D
6 2024 DigiDogs/D-Pose I SF [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Dogs PCK for 2D & 3D,
MPJPE for 3D
AMR
(Model-Based):
7 2017 SMAL I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} Quadruped Scan-to-Mesh Distance,
Keypoint and Silhouette
Reprojection Err.
8 2019 Three-D Safari I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Zebra PCK, IoU, Geodesic Distance
9 2020 SMBLD I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} StanfordExtra (2020) IoU, PCK
10 2021a Li et al. I S [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]} [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} BADJA (2018), PCK, IoU
Animal Pose (2019),
StanfordExtra (2020)
11 2021 Birds of a Feather I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Birds, PCK, IoU
StanfordExtra (2020)
12 2022 BARC I S [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} StanfordExtra (2020) PCK & IoU for 2D,
@@
13 2023 BITE I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} StanfordExtra (2020), Scan-to-Mesh Distance,
Semi-Synthetic 3D (2023) Mesh-to-Scan Distance,
@@
AMR
(Model-Free):
14 2022 BANMo I,V SF [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]} , [Uncaptioned image], {[Uncaptioned image], [Uncaptioned image]} Animals & Humans F-score,
3D Chamfer Distance (cm)
15 2023 DOVE I,V U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Bird, Horse, mIoU,
@@
16 2023 MagicPony I U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] COCO (2014), PCK,
Horse-10 (2021), 3D Chamfer Distance (cm)
DOVE (2023), others
17 2024b 3D-Fauna I U [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image] Quadruped Animals PCK@0.1, KT-PCK@0.1
3D APE from Image & Video - Multi-Camera-Views Methods
18 2019 Monet I SM [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Dog, Monkey PCK, Reprojection Err.
19 2020 OpenMonkeyStudio I S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Monkey Median & Mean Err. (cm),
Std (cm)
20 2021 GIMBAL V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mouse MPE (mm), RPA-MPE (mm)
21 2021 DAN-NCE V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Rats, Marmosets Mice, PCK, Reconstruction Err.
@@
22 2021 Anipose V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Fly, Mouse Joint Position Err. (mm)
23 2022 Muramatsu et al. V S [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] AcinoSet (2021) RMSE, PCK, $l^{2}$ norm in pixels
24 2023 3D-UPPER V U [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Mice RMSE, Scalar Product,
@@
25 2023 MAMMAL V S {[Uncaptioned image], [Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image], [Uncaptioned image] Pigs, Mouse, Dogs MPJPE, IoU,
@@
26 2024 SBeA V SF [Uncaptioned image], {[Uncaptioned image]+[Uncaptioned image]}, [Uncaptioned image], [Uncaptioned image] Mice, Dogs, Birds RMSE (pixels), Reprojection Err.
4.3.1 Single Camera View